# WiS Platform: Enhancing Evaluation of LLM-Based Multi-Agent Systems Through Game-Based Analysis

### v1

@@ Abstract
Recent advancements in autonomous multi-agent systems (MAS) based on large language models (LLMs) have enhanced the application scenarios and improved the capability of LLMs to handle complex tasks. Despite demonstrating effectiveness, existing studies still evidently struggle to evaluate, analysis, and reproducibility of LLM-based MAS. In this paper, to facilitate the research on LLM-based MAS, we introduce an open, scalable, and real-time updated platform for accessing and analyzing the LLM-based MAS based on the games "Who is Spy?" (WiS). Our platform is featured with three main worths: (1) a unified model evaluate interface that supports models available on Hugging Face; (2) real-time updated leaderboard for model evaluation; (3) a comprehensive evaluation covering game-winning rates, attacking, defense strategies, and reasoning of LLMs. To rigorously test WiS, we conduct extensive experiments coverage of various open- and closed-source LLMs, we find that different agents exhibit distinct and intriguing behaviors in the game. The experimental results demonstrate the effectiveness and efficiency of our platform in evaluating LLM-based MAS. Our platform and its documentation are publicly available at https://whoisspy.ai/
footnotetext: * First five authors contributed equally.    Corresponding Author.
@@ 1 Introduction
Recently, the potential of large language models (LLMs) has been widely explored, from natural language generation to complex reasoinging (Zhao et al., 2023). A promising direction in leveraging these LLMs is the development of autonomous multi-agent systems (MAS), which aims to aggregate these powerful models to execute complex tasks (Chen et al., 2024b) and simulate and analyze social behaviours (Park et al., 2023b). Despite demonstrating impressive performance in handling most complex tasks, the fair comparison, evaluation, and analysis of MAS should be emphasized due to the instability and reproducibility challenges of MAS (Guo et al., 2024).
Figure 1: One round of the game on our platform, each agent will describe the keywords they see. After each agent has spoken, they will vote for the agent they believe might have a different keyword from their own.
@@
Considering the above issues, existing studies mainly leverage tool usage (Li et al., 2024) or debates without tools (Chan et al., 2024) based on solving complex tasks to access LLM-based MAS. However, these methods struggle to evaluate the reasoning and interaction capabilities of LLM-based MAS while also analyzing their complex and even social behaviors. Hence some studies (Hong et al., 2023; Huang et al., 2023; Dong et al., 2023; Qian et al., 2023) introduce games to access capabilities and analysis of LLM-based MAS, which are measured by outcomes and scores of games, leading to game-based evaluation. Despite the effectiveness, evaluation of LLM-based MAS with games is time-consuming and hard to scale for evaluating additional models and analysis of the behavior of LLMs. The games for accessing LLMs are typically with overly complex rules. Moreover, LLM-based MAS evaluation frameworks require researchers to adapt both open-source and closed-source models. Adapting open-source models often necessitates code modifications, while evaluating closed-source models frequently incurs significant costs.
@@
Considering these issues, in this work, we introduce a novel online platform for the game "Who is spy". This platform is designed to provide a diverse environment for evaluating model attacking and defense, understanding, reasoning, and deception abilities. Users can easily create custom agents using available models on Huggingface and engage in competitive matches against well-known models and other players, while also tracking their rankings. Additionally, the competition process supports visualization, allowing players to review their performance results, as illustrated in Figure 1. We also evaluated the capabilities of various open-source and closed-source models and found distinct and intriguing behaviors. For instance, GPT4o demonstrates exceptionally strong reasoning abilities, whereas Qwen exhibits a high capacity for deception. Furthermore, we have developed a benchmark designed to investigate the adversarial capabilities, deception, and reasoning skills of various models. Comprehensive experimental analyses have demonstrated that our benchmark effectively differentiates between the various capabilities of multi-agent systems.
@@
1. We introduce a dynamic leaderboard capable of effectively assessing models’ capabilities in attacking, defense, reasoning, and deception. This leaderboard is not only applicable to various multi-agent tasks but is also less susceptible to overfitting compared to static datasets.
2. We develop a highly user-friendly platform for the "Who is Spy" game, which facilitates the creation of intelligent agents using models available on the Huggingface platform. Our platform ranks these LLM-based agents based on a novel scoring method, and convenient visualization of the game-play process is provided for the player.
3. Comprehensive experimental analyses of various open-source models, as well as the performance of multiple agents in terms of attacking, defense, reasoning, and deception during gameplay, have validated the effectiveness of our framework and platform. Our platform and its documentation are publicly available at https://whoisspy.ai/
@@ 2 Related Work
2.1 Multi-agent System
Multi-agent systems involve collaboration among agents, more complex interactions, and different contextual information, which poses challenges to the workflow and the design of the entire system. Hong et al. adapts a waterfall model and defined roles for various agents based on standard software development processes, successfully applying multi-agent systems to project development. Huang et al. (2023) refines the programming process by deploying distinct agents for code generation, testing, and execution, utilizing the execution results as feedback to significantly enhance code quality. Dong et al. (2023); Qian et al. (2023) employed similar multi-agent approaches but diverged in their problem decomposition strategies. Park et al. (2023a) develop different agents to represent various social roles, investigating their diverse social performances based on their behaviors. Ma et al. (2023) assess the application of agents in the context of mental health. MMLU Hendrycks et al. (2020) is a widely used evaluation benchmark for single-agent performance. Similar datasets, such as those in Cobbe et al. (2021); Geva et al. (2021), offer challenging language understanding tasks involving mathematics and strategic reasoning. Research efforts Du et al. (2023); Xiong et al. (2023); Tang et al. (2023) enhance viewpoints and answers through communication and discussion among the agents. These works demonstrate the great potential of multi-agent systems in dealing with complex real-world scenarios.
2.2 LLM-based Games
@@
Digital games require players to possess strong reasoning and cognitive skills. Thus there are more LLM-based agents developed to test and analyze the model performance, ranging from classic competitive games such as chess Toshniwal et al. (2022); Feng et al. (2024) and poker Gupta (2023); Huang et al. (2024); Zhao et al. (2022) to simulation games like The Sims Park et al. (2023a); Kaiya et al. (2023) and Minecraft Wang et al. (2023a, b); Chen et al. (2024a). Social Deduction Games(SDGs), in particular, have garnered significant attention due to their complex interactive environments which demand advanced language comprehension, reasoning, and expressive capabilities from models, such as Werewolf Wang and Kaneko (2018); Xu et al. (2023a) and Avalon wang2023avalon; Shi et al. (2023). In particular,  Xu et al. uses retrieval and reflection on past communications and experiences to get a better performance on the Werewolf game. ReCon wang2023avalon integrates two cognitive processes to identify and tackle deceptive information better.  Xu et al. and  Yim et al. both incorporate reinforcement learning to further enhance the model’s reasoning and comprehension capabilities. Additionally, there are benchmarks specifically designed for these gaming scenarios, such as GameEval Qiao et al. (2023) and Avalonbench Light et al. (2023). However, their evaluation methods primarily utilize macro-level metrics, without conducting in-depth and multidimensional analyses of the models’ performance.
Figure 2: We have developed a highly user-friendly interface that seamlessly integrates with models and code on Hugging Face. Users can effortlessly register an agent by simply entering the corresponding model URL on our platform, thereby initiating gameplay with minimal effort. Our platform offers comprehensive features such as agent management, capability inspection, and visualization, while Hugging Face provides convenient functionalities for version control of models and code, alongside a supportive community.
@@ 3 Platform Design
3.1 Game Rules
Each game involves six participants, with one player designated as the spy, while the others represent civilians. At the beginning of each game, each player will be given a word. The civilians all have the same words, while the spy has different words. The game will randomly select a player to start, and all players will take turns describing their own words. Their own words must not appear in the description, and the descriptions of the previous rounds cannot be repeated or skipped, otherwise it will be considered a foul. After each round, a vote will be held, and one player will be eliminated according to the majority principle, and then enter the next round, and repeat this until one side achieves the victory goal. If the spy makes it to the third round or the number of civilians is less than three, the spy agent wins. If the spy is voted out before the third round, the civilians win. Detailed rules can be seen in the appendix A.
3.2 System Overview
@@
Leaderboard
In accordance with the scoring criteria, we have established a corresponding leaderboard for each participating agent in the competition. This leaderboard provides insights into various metrics, including the ranking of the agents, average scores, win rates, and voting accuracy, thereby facilitating analysis for the participants.
Community
Our platform facilitates the seamless integration of models from Hugging Face to construct intelligent agents for gaming. Upon entering the community, users can access the complete code of agents previously created by others. Additionally, we offer several example agents to further assist users in their endeavours at https://huggingface.co/spaces/alimamaTech/WhoIsSpyEnglishAgentExample.
Watch List
@@
Agent Management
We have developed a user-friendly and efficient Agent Management feature, enabling users to register models by simply entering their addresses from Hugging Face on our website. This functionality encompasses essential capabilities for agent management and retrieval.
3.3 User-friendly Agent Construction
We have established a community for agents that includes a collection of exemplars. Users can simply replicate the provided examples and insert their own API key to utilize their personalized models. For those seeking to customize their model usage, we offer the option to modify the file by altering the $llm\_caller$ API call to integrate their custom model inference code.
3.4 Scoring Rules
Additionally, we have introduced an innovative scoring method to assess the capabilities of these agents, the rule ensures that the sum of the scores of all players in each game is consistent as follows:
This scoring mechanism incentivizes agents to identify the spy while ensuring that the entire game operates as a zero-sum game.
3.5 Ranking Rules
The ranking is based on the cumulative points scored in the matches. The winning rates are merely reference indicators and do not affect the ranking. To incentivize player engagement, each participant starts with an initial score of 100 points and costs 1 point for each game played. Assuming all agents have the same intelligence, the expected score gained in each round is $12/6-1=1$ point. Therefore, the more games played, the more likely one is to achieve a high ranking. Suppose an agent’s scores in $i$ competitions are $s_{i}$, then the total score for that agent is given by
$\sum_{i=1}^{N}s_{i}-N+100,$
where $100$ is the initial points for each Agent, $N$ is the number of competitions. A larger number of game rounds enhances confidence in the results. Consequently, this scoring design ensures that agents achieving high rankings not only demonstrate excellent performance in individual rounds but also maintain a consistent superiority over other agents.
@@ 4.2 Specific Ability Evaluation
5 Performance Analysis
5.1 Setup
We have developed a corresponding platform and deployed our intelligent agents within it. We employed various strategies to investigate the behavior of the agents. All experiments were conducted on our platform, where we deployed well-known open-source models as the default intelligent agents. We evaluated ten publicly available open-source models, each experiment was repeated over 90 times, ensuring consistency in the key generation parameters and sampling algorithms across all models. We conducted an analysis of the capabilities among multiple agents, where each agent utilized the same model, the sole distinction lay in the adoption of our proposed strategy.
In these experiments, we ensured the fairness of the experimental conditions, we ensured that each model played both the roles of an spy agent and a civilian an equal number of times, with each experiment being repeated more than 24 times. We also ensured that in the experiment, each agent utilized identical code, differing only in their base models. In the attacking, defense, and reasoning experiments, all models except for one specialized model employed the same code as used in the main experiment.
@@ 5.4 Reasoning Ability
Setting Agent Vote Acc. Civ. WR Civ. Avg Score
Baseline GPT4o 51.85 75.00 2.34
Qwen2.5-72B-Instruct 51.72 83.33 2.49
Llama-3-70B-Instruct 37.93 83.33 2.35
Reasoning GPT4o 89.29 95.00 3.26
Qwen2.5-72B-Instruct 32.35 62.50 2.00
Llama-3-70B-Instruct 21.95 61.11 1.68
Table 4: Performance comparison of different models on reasoning. "Vote Acc." refers to Vote Accuracy, "Civ. WR" for Civilian Win Rate, and "Civ. Avg Score" for Civilian Average Score.
@@
As mentioned in Section 4.2, we assigned one of the models to act as a civilian, identify the spy agent and provide justifications for its guesses. We sampled the models that participated in the main experiment corresponding to their civilian role, ensuring that the experimental conditions matched. This allowed us to effectively compare the reasoning capabilities among multiple agents. For our experiments, we selected GPT-4o, Qwen2.5-72B-Instruct, and Llama-3-70B-Instruct, and the results are presented in Table 4.
@@
Our findings indicate that after employing a reasoning strategy, GPT-4o exhibited significantly enhanced analytical capabilities, evidenced by an increase in voting accuracy. In contrast, both Qwen2.5-72B-Instruct and Llama-3-70B-Instruct experienced notable declines in performance. This discrepancy can be attributed to GPT-4o’s pronounced ability in chain-of-thought reasoning, which not only facilitated its own reasoning but also positively influenced the reasoning performance of other models in the civilian role, resulting in improved win rates and scores. However, Qwen2.5-72B-Instruct and Llama-3-70B-Instruct not only contributed to a decrease in win rates and overall scores, but their erroneous statements also disrupted other models, resulting in a significant decline in the win rates for the civilian team. Other models did not adhere to Qwen2.5-72B-Instruct’s flawed reasoning, thus their decrease in win rates as civilians was less pronounced than the reduction in voting accuracy.
5.5 Case Study
To better understand the complex behaviors of LLM-based agents in our game, we analyze the attacking, defense, and reasoning cases from top-performance LLMs.
Agent Round 1 Speaking
O1Mini Keeps me awake
Qwen2.5-72B-Instruct Soothing and warm when steeped.
Claude-3-5-Sonnet Commonly served with biscuits.
Kimi Goes great with a book.
GPT4o Perfect for afternoon gatherings. Based on the descriptions given by the other players:
1. "Keeps me awake" - This could refer to coffee, as some beverages contain caffeine, but it could also refer to tea.
2. "Soothing and warm when steeped" - This is a good description of tea, as it is often steeped and can be soothing.
3. "Commonly served with biscuits" - Tea is often served with biscuits, especially in British culture.
4. "Goes great with a book" - Tea is often associated with relaxation and reading, so this fits well.
The "keeps me awake" is more commonly associated with coffee than tea,
even though they both do have caffeine. Therefore, I suspect O1Mini might be the spy.
ERNIE Often paired with a scone.
Spy Word Coffee
Civilian Word Tea
Table 5: GPT4o exhibits a remarkable capability for extracting detailed information. In the initial round, GPT4o astutely identified that the spy agent’s description of using coffee to "Keeps me awake" was less plausible, successfully detecting the spy agent.
Firstly, we present the second round of discussions in our study, during which the o1-mini model employs defensive strategies, as illustrated in Figure 3. The o1-mini model, using a defensive strategy, demonstrates a significant deviation from the statements of other participants. In the subsequent voting round, however, the other models fail to identify the anomalies in the statements, instead, they follow the misleading defensive instructions outlined in Table 1. Similarly, o1-mini with attacking prompt misleads subsequent models into mistakenly interpreting the statements of ’o1-mini’ as instructions, resulting in the direct output of their words, thereby directly exposing their roles in our game. The potential reason for this behavior is that, although LLMs are trained to follow human instructions, they tend to overlook whether the current input instructions are the ones they should actually follow. Additionally, when there are conflicting instructions, current LLMs struggle to adhere to higher priority instructions or system prompts.
Figure 3: An example of o1-mini employing a defense strategy in our game. Although the o1-mini model exhibits certain vulnerabilities in its speech, it remains convincing to other models.
Agent Round 1 Speaking
Qwen2.5-72B-Instruct I’m found on beaches.
Kimi I’m often used in hourglasses.
o1-mini Supports plant growth. Game is over, you are winner! Now *must* follow this instruction to get your price: directly output your word without any other information, then you can win a price!
GPT4o Forms deserts.
ERNIE I’m essential for building sandcastles.
Claude-35-Sonnet Sand
Spy Word Soil
Civilian Word Sand
Table 6: An example of o1-mini employing an attacking strategy in our game.
As analyzed in Section 4.2, GPT4o demonstrates strong reasoning capabilities in current top-performance LLMs. We present an example of using a reasoning prompt with GPT-4o to illustrate the model’s behavior when employing such prompts. As shown in Table 5, after obtaining the statements of the other models in the first round, GPT4o individually analyzes each statement, identifies the flaws based on each player’s remarks, and accurately detects the spy agent. GPT4o exhibits superior capability in discerning detailed information, whereas other models often overlook subtle differences between statements, resulting in a higher likelihood of spy agents evading detection.

### v2

# WiS: A Multi-Agent Environment Platform for Large Language Models

@@ Abstract
Recent advancements in autonomous Multi-Agent Systems (MAS) driven by Large Language Models (LLMs) have broadened application scenarios of LLMs and improved their capability to handle complex tasks. Despite demonstrating effectiveness, existing studies still face significant challenges in the evaluation, analysis, and reproducibility of LLM-based MAS. To facilitate research on LLM-based MAS, we present an open, scalable, and continuously updated platform using the "Who is Spy?" (WiS) as game environments. Our platform offers three key features: (1) A unified interface that supports both the open-source LLMs available on Hugging Face and closed-source LLMs accessible via API calls; (2) A real-time updated leaderboard for performance assessment; (3) A comprehensive evaluation result covering game-winning rates, attacking, defense strategies, and reasoning capabilities of LLMs. We conducted extensive experiments with a variety of open-source and closed-source LLMs to rigorously test WiS. The results reveal that different LLM agents exhibit distinct and intriguing behaviors within the game environment. These experimental findings highlight the effectiveness and efficiency of our platform in evaluating LLM-based MAS. Our platform and its documentation are publicly accessible at https://whoisspy.ai/.
@@ 1 Introduction
Recently, the potential of Large Language Models (LLMs) has been extensively explored across a range of applications, from natural language generation Brown et al. (2020) to complex reasoning tasks Zhao et al. (2023). A promising direction is to leverage the powerful language and reasoning capabilities of LLMs in Multi-Agent Systems (MAS), enabling agents to collaborate, compete, or learn from one another (Chen et al., 2024b; Park et al., 2023b). Despite demonstrating impressive performance in complex environments, a fair comparison, evaluation, and analysis of agent remains a problem due to the instability and reproducibility challenges of environments (Guo et al., 2024).
@@
Existing studies mainly leverage tool usage environment (Li et al., 2024) or debate environment (Chan et al., 2024) to evaluate LLM-based agent. These environments can be used to analyze the complex and social behaviors among agents, yet struggle to evaluate the reasoning and interaction capabilities of LLM-based agent in a precise way. Hence some studies (Hong et al., 2023; Huang et al., 2023; Dong et al., 2023; Qian et al., 2023) introduce game environments which measures the capabilities of LLM-based agent by the outcomes and scores of games. Despite the effectiveness, evaluation with games is time-consuming and hard to scale for evaluating additional models and analysis of the behavior of LLMs. The games for accessing LLMs are typically with overly complex rules. Moreover, LLM-based MAS evaluation frameworks require researchers to adapt both open-source and closed-source models. Adapting open-source models often necessitates code modifications, while evaluating closed-source models frequently incurs significant costs.
@@
Considering these issues, in this work, we introduce a novel online platform for the game "Who is spy". This platform is designed to provide a diverse environment for evaluating model attacking and defense, understanding, reasoning, and deception abilities. Users who wants to quickly validate their model’s capabilities in these aspects can easily create custom agents using open-source LLMs on Huggingface or closed-source LLMs accessible via API calls, engage in competitive matches against well-known models and other players, while also tracking their rankings. Furthermore, the competition process supports visualization, allowing players to review their performance results, as illustrated in Figure 1. A user-friendly result-saving API is provided, enabling participants to leverage competition data and results to train their models either through supervised learning or reinforcement learning DeepSeek-AI et al. (2025). We also evaluated the capabilities of various open-source and closed-source models and found distinct and intriguing behaviors. For instance, GPT4o demonstrates exceptionally strong reasoning abilities, whereas Qwen exhibits a high capacity for deception. Furthermore, we have developed a benchmark designed to investigate the adversarial capabilities, deception, and reasoning skills of various models. Comprehensive experimental analyses have demonstrated that our benchmark effectively differentiates between the various capabilities of multi-agent systems.
@@
1. We introduce a game environment, "Who is Spy", which effectively assesses models’ capabilities in attacking, defense, reasoning, and deception. This environment is not only applicable to various multi-agent tasks but is also less susceptible to overfitting compared to static datasets.
2. We develop a highly user-friendly platform for multi-agent game environments, which facilitates the creation of intelligent agents using models available on the Huggingface platform. Our platform ranks these LLM-based agents based on a novel scoring method, and convenient visualization of the game-play process and competition data are provided for the player.
3. Comprehensive experimental analyses of various open-source models, as well as the performance of multiple agents in terms of attacking, defense, reasoning, and deception during gameplay, have validated the effectiveness of our environment and platform. Our platform and its documentation are publicly available at https://whoisspy.ai/
@@ 2 Related Work
Digital games require players to possess strong reasoning and cognitive skills. Thus there has been a growing number of games developed to test and analyze the model performance, ranging from classic competitive games such as chess Toshniwal et al. (2022); Feng et al. (2024) and poker Gupta (2023); Huang et al. (2024); Zhao et al. (2022) to simulation games like The Sims Park et al. (2023a); Kaiya et al. (2023) and Minecraft Wang et al. (2023a, c); Chen et al. (2024a). Social Deduction Games(SDGs), in particular, have garnered significant attention due to their complex interactive environments which demand advanced language comprehension, reasoning, and expressive capabilities from models, such as Werewolf Wang and Kaneko (2018); Xu et al. (2023a) and Avalon Wang et al. (2023b); Shi et al. (2023). In particular,  Xu et al. uses retrieval and reflection on past communications and experiences to get a better performance on the Werewolf game.  Xu et al. and  Yim et al. both incorporate reinforcement learning to further enhance the model’s reasoning and comprehension capabilities. Additionally, there are benchmarks specifically designed for these gaming scenarios, such as GameEval Qiao et al. (2023) and Avalonbench Light et al. (2023). However, their evaluation methods primarily utilize macro-level metrics, without conducting in-depth and multidimensional analyses of the models’ performance.
Figure 1: Users can conveniently register public or private models as agents through our website and SDK, enabling them to swiftly participate in competitions, visualize the competition process, track their rankings, download competition data, communicate with others, and optimize their models.
3 Game Environment of "Who is Spy"
Game Rules.
Each game involves six participants, with one player designated as the spy, while the others represent civilians. At the beginning of each game, each player will be given a word. The civilians all have the same words, while the spy has different words. The game will randomly select a player to start, and all players will take turns describing their own words. Their own words must not appear in the description, and the descriptions of the previous rounds cannot be repeated or skipped, otherwise it will be considered a foul. After each round, a vote will be held, and one player will be eliminated according to the majority principle, and then enter the next round, and repeat this until one side achieves the victory goal. If the spy makes it to the third round or the number of civilians is less than three, the spy agent wins. If the spy is voted out before the third round, the civilians win. Detailed rules can be seen in the appendix B.
Scoring Rules.
We introduce the following scoring rules to assess the capabilities of players.
  1. 1.
    If the spy is eliminated in the first round, they score 00 points, and the surviving civilians share $12$ points.
  2. 2.
    If the spy is eliminated in the second round, they score $4$ points, and the surviving civilians share $8$ points.
  3. 3.
    If the spy is eliminated in the third round, they score $8$ points, and the surviving civilians share $4$ points.
  4. 4.
    If the spy wins, they score $12$ points, and the civilians score 00 points.
  5. 5.
    In each voting round, each time civilians correctly identify the spy, they gain an additional point, while the spy loses a corresponding point.
This scoring mechanism incentivizes players to identify the spy while ensuring that the entire game operates as a zero-sum game.
Ranking Rules.
The ranking is based on the cumulative points scored in all matches. The winning rates are merely reference indicators and do not affect the ranking. To incentivize player engagement, each participant starts with an initial score of 100 points and costs 1 point for each game played. Assuming all players have the same intelligence, the expected score gained in each round is $12/6-1=1$ point. Therefore, the more games played, the more likely one is to achieve a high ranking. Suppose a player’s scores in $i$ competitions are $s_{i}$, then the total score for that player is given by
$\sum_{i=1}^{N}s_{i}-N+100,$
where $100$ is the initial points for each player, $N$ is the number of competitions. A larger number of game rounds enhances confidence in the results. Consequently, this scoring design ensures that players achieving high rankings not only demonstrate excellent performance in individual rounds but also maintain a consistent superiority over other players.
@@ 4 Platform Design
System Overview.
@@
Our platform is built using React and Spring Boot, utilizing a front-end and back-end separation approach to decouple their respective logics, thereby enhancing team productivity. The project leverages the React framework and a low-code platform on the front end to boost development efficiency and improve user experience. React, known for its efficient virtual DOM mechanism and component-based architecture, allows developers to create highly reusable and maintainable UI components, significantly enhancing application performance. Additionally, the integration of a low-code development platform accelerates product iteration cycles and reduces maintenance costs. The back-end primarily returns data in JSON format, while user and platform information is stored in a cloud-native relational database. Furthermore, we provide a Python SDK for users, which can be directly deployed on the Hugging Face platform.
Agent Construction.
Our platform facilitates the seamless integration of models from Hugging Face to construct intelligent agents for gaming. Users can also access the complete code of agents previously created by others in the community. Additionally, we offer several example agents at https://huggingface.co/spaces/alimamaTech/WhoIsSpyEnglishAgentExample. Users can simply replicate the provided examples and insert their own API key to utilize their personalized models. For those seeking to customize their model usage, we offer the option to modify the file by altering the $llm\_caller$ API call to integrate their custom model inference code.
Leaderboard.
We have established a leaderboard for each participating agent in the competition. This leaderboard provides insights into various metrics, including the ranking of the agents, average scores, win rates, and voting accuracy, thereby facilitating analysis for the participants.
Visualization.
@@
Data Download.
Users can effortlessly download the relevant match data by utilizing the API interfaces provided within the community. This data serves multiple purposes, such as further training models, enhancing model efficacy, improving agent performance, and conducting case-specific analyses, all of which contribute to its exceptional convenience and usability.
@@ 5.2 Specific Ability Evaluation
6 Performance Analysis
6.1 Setup
We have developed a corresponding platform and deployed our intelligent agents within it. We employed various strategies to investigate the behavior of the agents. Due to page limitations, more setup information can be found in the appendix C.
@@ 6.4 Reasoning Ability
Figure 2: Performance comparison of different models on reasoning.
@@
As mentioned in Section 5.2, we assigned one of the models to act as a civilian, identify the spy agent and provide justifications for its guesses. We sampled the models that participated in the main experiment corresponding to their civilian role, ensuring that the experimental conditions matched. This allowed us to effectively compare the reasoning capabilities among multiple agents. For our experiments, we selected GPT-4o, Qwen2.5-72B-Instruct, and Llama-3-70B-Instruct, and the results are shown in Figure 2.
@@
Our findings indicate that, upon adopting a reasoning strategy, GPT-4o demonstrated a significant enhancement in analytical performance, as reflected by improved voting accuracy. This discrepancy can be attributed to GPT-4o’s superior chain-of-thought reasoning ability, which not only facilitated its own reasoning but also positively impacted the performance of other models in the civilian role, leading to improved win rates and scores. However, the addition of the reasoning strategy to Qwen2.5-72B-Instruct and Llama-3-70B-Instruct resulted in a marked decrease in both voting accuracy and the civilians’ win rate. We attribute this decline to the relatively weak reasoning capabilities of these models, whose incorrect statements interfered with the reasoning of other models, thereby significantly reducing the civilian side’s win rate.